# Distributed AI Platform for the 6G RAN

### v1

@@ VII Open vs Closed architecture and interfaces for integration with RAN
Despite the progress within the O-RAN community, many fundamental tensions remain around control interfaces. For example, several major vendors do not support the idea of a near real-time RIC [15], expressing concerns regarding the network stability and security. If RAN vendors allow developers to exert fine-grained control on the RAN behavior, these may clash with the proprietary algorithms vendors have implemented. Also, some vendors argue that exposing sensitive RAN data may affect the security of the network and the data privacy. Vendors are also reluctant to open up some of their interfaces, as this may reduce their competitive advantage. All these concerns are important, and different operators take different stances on them. Our view is that the distributed AI platform proposed in this article should be a building block that can be customized appropriately for different use cases. Next, we briefly discuss some examples of deployment options (also illustrated in Fig. 4).
Figure 6: Deployment options for integration of distributed AI platform with the RAN.
@@
Proprietary design – Another extreme is a proprietary design, fully controlled by the RAN vendor (Fig. 6, middle). In this case, the group that is in charge of the RAN product development could also be in charge of developing and managing the AI platform and of defining proprietary interfaces for data collection and control. A different group of the RAN vendor could deploy and manage their own AI RAN applications, without having to constantly go through the RAN product development group, to ask them to introduce new features or expose more data. By decoupling the responsibilities, the vendor can accelerate its innovation, while still maintaining a full control over the RAN behavior.
Hybrid design – It is also possible to create a hybrid solution (Fig. 6, right), that would be a mix of proprietary AI processor runtimes and runtimes running on top of standard O-RAN RICs (e.g., only non RT-RIC). AI applications could be composed of a mixture of first-party and third-party AI blocks, where the proprietary data and control interfaces of the RAN functions can only be accessed by the first-party AI blocks, whereas thrid-party blocks can only access APIs exposed by the standard interfaces. This would allow the RAN vendor to maintain control over the RAN behavior, while still allowing third-party developers to innovate.
VIII Conclusions
In this paper we focused on the topic of AI as a key enabler behind realizing the 6G vision. We argued about the benefits of introducing AI across several dimensions of the RAN, from the management and infrastructure up to the application layer. Based on these observations, we outlined the challenges for deploying AI solutions, stemming from the requirements of the applications and the characteristics of the RAN infrastructure. Motivated by these challenges, we proposed a distributed AI platform architecture. Its goal is to alleviate the common painpoints of deploying AI applications in the RAN without being prescriptive about the implementation details. This allows the platform to be tailored to the needs of the infrastructure provider or a standardization body. We believe that by identifying the future requirements and by initiating a discussion on the architecture of a 6G AI platform we can help both the standards and the vendors to create new opportunities for introducing AI solutions in this space.
Ganesh Ananthanarayanan is a Senior Principal Researcher at Microsoft. His research focuses on systems and networking, with current work on AI inference systems. He received his PhD from UC Berkeley.
Xenofon Foukas is a Principal Researcher at Microsoft. His research interests are on networks and distributed systems with his current focus on edge computing. He received his PhD from the University of Edinburgh.
Božidar Radunović is a Senior Principal Researcher at Microsoft Research. His research interests are in design and building next generation cloud and edge infrastructure. He received his PhD from EPFL.
Yongguang Zhang is currently a Partner Principal Researcher at Microsoft Research. He is currently leading the cloud RAN platform R&D. He received his Ph.D. in Computer Science from Purdue University.

### v2

@@ 7. Open vs Closed architecture and interfaces for integration with RAN
Despite the progress within the O-RAN community, many fundamental tensions remain around control interfaces. Several major vendors do not support the idea of a near-RT RIC (Morris, 2024), expressing concerns regarding the network stability and security. If RAN vendors allow developers to exert fine-grained control on the RAN behavior, these may clash with the proprietary algorithms vendors have implemented. Also, some vendors argue that exposing sensitive RAN data may affect the security of the network and the data privacy. Vendors are also reluctant to open up some of their interfaces, as this may reduce their competitive advantage. Our view is that the distributed AI platform proposed in this work should be a building block that can be customized appropriately for different use cases. We briefly discuss some examples of deployment options (Fig. 4).
Figure 8. Deployment options for integration of distributed AI platform with the RAN.
@@
Proprietary design – Another extreme is a design, fully controlled by the RAN vendor (Fig. 8, middle). In this case, the group that is in charge of the RAN product development could also be in charge of developing and managing the AI platform and of defining proprietary interfaces for data collection and control. A different group could deploy and manage their own AI RAN applications, without having to constantly go through the RAN product development group, to ask them to introduce new features or expose more data. By decoupling the responsibilities, the vendor can accelerate its innovation, while still maintaining a full control over the RAN behavior.
Hybrid design – It is also possible to create a hybrid (Fig. 8, right), that would be a mix of proprietary AI processor runtimes and runtimes running on top of standard O-RAN RICs. AI applications could be composed of a mixture of first-party and third-party AI blocks, where the proprietary data and control interfaces of the RAN functions can only be accessed by the first-party AI blocks, whereas third-party blocks can only access APIs exposed by the standard interfaces. This would allow the RAN vendor to maintain control over the RAN behavior, while still allowing third-party developers to innovate.
8. Conclusions
To realize the 6G vision, we argued about the benefits of introducing AI in the RAN, from the management and infrastructure up to the application layer. Based on these observations, we outlined the challenges for deploying AI solutions, stemming from the requirements of the applications and the characteristics of the RAN infrastructure. Motivated by these challenges, we proposed a distributed AI platform architecture. Its goal is to alleviate the common painpoints of deploying AI applications in the RAN without being prescriptive about the implementation details. By identifying the future requirements and by initiating a discussion on the architecture of a 6G AI platform will help both the standards and the vendors to create opportunities for introducing AI solutions in 6G.